A jazz group on tour has been drawing average crowds of 500 persons. It is projected that for every $1 increase in the $12 ticket price, the average attendance will decrease by 50. At what ticket price will nightly receipts be $5600?

Answer by richwmiller(17219)

12x=500 12*500=6000
13x=450 13*450=5850
14x=400 14*400=5600
(12+x)*(500-50x)=y
(12+x)*(500-50x)=5600
x=12
12+2=14 cost of ticket
500-50*2=400
14*400=5600